OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y = x + 2. A) y = 4x - 6 B) y = -4x - 6 C) y = -4x + 2 D) y = 4x

OpenStudy (ghazi):

since the line is perpendicular to the give line therefore slope of required line will be negative of inverse of given line's slope that is -1 now your equation of required line is y=-1*x+d...to find d substitute the given points (-1,-2)..you'll get d

OpenStudy (anonymous):

Okay thank you:) !!! @ghazi

OpenStudy (ghazi):

or else you can write equation directly as y-(-2)= -1(x-(-1))